Use the square root property to solve the equation. (x-6)^2=4
Found 2 solutions by algebrahouse.com, Edwin McCravy:
Answer by algebrahouse.com(1659)

(x - 6)² = 4
x - 6 = 2 or x - 6 = -2 {took the square root of both sides}
x = 8 or x = 4 {added 6 to both sides}

For more help from me, visit: www.algebrahouse.com



Answer by Edwin McCravy(20054)

(x - 6)² = 4

Take away the square on the left, take ± the square root of the
right side:
            _
  x - 6 = ±Ö4

  x - 6 = ±2

      x = 6 ± 2

Using the +        Using the -

x = 6 + 2          x = 6 - 2
x = 8              x = 4

Edwin